OpenStudy (anonymous):

tan20 + tan 40 + tan 60 +....tan180

OpenStudy (raden):

hint : tan180 = 0 tan160 = tan(180-20) = - tan20 tan140 = tan(180-40) = - tan40 tan120 = tan(180-60) = - tan60 tan100 = tan(180-80) = - tan80

OpenStudy (anonymous):

THANKS

OpenStudy (raden):

well, if u subtitute the values in my hint their sum be tan20+tan40+tan60+tan80+tan100+tan120+tan140+tan160+tan180 = tan20+tan40+tan60+tan80-tan80-tan60-tan40-tan20+0 = (tan20-tan20)+(tan40-tan40)+(tan60-tan60)+(tan80-tan80)+0 = 0